Evaluate the derivative lim┬(x→1)⁡[1/(1-x) ∫_(x^2)^x▒(t^2-1)dt]


1
Expert's answer
2021-04-13T12:24:50-0400

"lim_{x\rightarrow1}[\dfrac{1}{1-x}\int_{-x^2}^{x^2}(t^2-1)dt]"


"=lim_{x\rightarrow1}[\dfrac{1}{1-x}(\dfrac{t^3}{3}-t)|_{-x^2}^{x^2}"


"=lim_{x\rightarrow1}[\dfrac{1}{1-x}(\dfrac{x^6}{3}-x^2+\dfrac{x^6}{3}-x^2)"


"=lim_{x\rightarrow1}[\dfrac{1}{1-x}2x^2(\dfrac{x^3}{3}-1)"


"=\dfrac{2(1)^2(\dfrac{1}{3}-1)}{1-1}=\infty"


Hence The given limit does not exist.